Problem 5-10 (Algo)

The following represents a project that should be scheduled using CPM:

IMMEDIATE PREDECESSORS TIMES (DAYS)
ACTIVITY a m b
A 1 4 7
B 1 3 5
C A 2 5 11
D A 1 8 9
E B 1 2 3
F C,D 1 6 11
G D,E 1 2 3
H F,G 2 3 3


b. What is the critical path?

a. A-D-F-H

b. B-E-G-H

c.A-D-G-H

d.A-C-F-H

c. What is the expected project completion time? (Do not round intermediate calculations. Round your answer to 3 decimal places.)

d. What is the probability of completing this project within 22 days? (Use Excel's NORM.S.DIST() function to find the correct probability for your computed Z-value. Do not round intermediate calculations. Round "z" value to 2 decimal places and final answer to 4 decimal places.)
Answer #1
Activity Optimistic time-a Expected completion time-m Pessimistic time-b Expected time= (a+4*m+ b)/6 Variance, (sigma)^2= (b-a/6)^2
A 1 4 7 4.00 1.00
B 1 3 5 3.00 0.44
C 2 5 11 5.50 2.25
D 1 8 9 7.00 1.78
E 1 2 3 2.00 0.11
F 1 6 11 6.00 2.78
G 1 2 3 2.00 0.11
H 2 3 3 2.83 0.03

PATHS

Path Duration Variance
ADFH 19.833 5.583
BEGH 9.833 0.667
ADGH 15.833 5.056
ACFH 18.333 1.889

Answer b: The critical path is the longest path i.e ADFH

Answer c: the expected project completion time through critical path is 19.833

Answer d: the probability of completing this project within 22 days is 0.8204

step 1 we will find the variance of the tasks which lie on critical path 5.58
step 2 mean project time (u) of critical path is= 19.83
step 3 Required completion time is 22
step 4 standard deviation= sqrt(variance)=   2.363
step 5 because Z= (given completion time- u)/standard deviation 0.917
step 6 P(z)= using NORM.S.DIST(z,true) 0.8204

Please ask, if you have any doubts through the comment section. Do rate the answer